OpenStudy (anonymous):

Find tan of angle B. Write your answer as a simplified fraction. Use the triangle below to answer this quesiton. Angle B= 30 degrees

OpenStudy (anonymous):

find the point on the unit circle corresponding to the 30 degree angle on the last page of the attached cheat sheet the first coordinate is cosine the second coordinate is sine you want \(\frac{\sin(30)}{\cos(30)}\) i.e. the second coordinate divided by the first

OpenStudy (anonymous):

did you find the point?

OpenStudy (anonymous):

btw if you have a triangle in front of you, and one angle is 30 degrees, take "opposite over adjacent" for your tangent.

OpenStudy (anonymous):

Okay I think I understand that question....it makes sense now that I can really see and understand what I'm doing....thanks!

OpenStudy (anonymous):

yw